UT Sep 08 2022 - Night #7 of Restart / Split with PEPSI-PETS




Night Info

  • Sunset 18:36
  • Nautical Twilight Ends 19:30
  • Astro Twilight Ends 20:00
  • Astro Twilight Begins 04:35
  • Nautical Twilight Begins 05:05
  • Sunrise 05:59
  • Moonset 03:53
  • Moon illumination ~93%
  • Handover to PEPSI/PETS 21:15 local time
  • PEPSI-PETS 21:15-01:18
  • Handover from PEPSI-PETS to Restart/Eng ~01:20

Staff

  • OSA: J.Williams
  • ISp: J.Schuch
  • Mtn Manager: T.Castro
  • ISA/SO: J.Power, B.Rothberg, R.T.Gatto
  • Technical Expert: J.Hill
  • AO: D. Miller
  • PEPSI: I.Ilyin

Night Plan

ALTA prediction for Tonight

20220907_see_evol_time_Height_20_20000.png

NOTE: Tonight's plans are highly variable and contingent upon several factors (LUCI SW fix, AO release, sky conditions, analysis of Field Aberration data or AGW Transfor data)

LUCI-1 MCE/ROE issues - may be offline

Starting Configuration LUCIs

Telescope Wakeup Continued

  • Conifrmatory Field Aberration (2 hrs) - Critical (1hour)
    • Requires preliminary analysis to be completed
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute - DOES NOT NEED LUCIs - Just AGW 1 and AGW 2
    • must complete SH spot centering and hotspot check before executing
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LUCI Checkout (if LUCI SW problem is fixed and Field Aberration Confirmatory data not requred)

  • Collect transform data - (2hrs) - confirmatory - Requires LUCIs
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
    • Can be done wth LUCI SW in crippled state (does not need flex comp, remember to load EmptyQueue.xml)

  • If Transform data not needed run these - can be run with LUCI-2 only if LUCI-1 MCE/ROE is offline (L1 and L2 versions exist for all scripts)
    • LUCI Scripts:
      • LUCI_2022_Restart_V1.80.xml - all scripts (except LS Rot ZP)
      • Rot_ZP_V2 script
      • LUCI_Phot_ZP_OT_V6.xm - LUCI photometric standards (fixes AGW patrol problem)
    • Imaging check (1hr)
      • ANY seeing OK
      • dither in-field
        • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am
      • dither out-of-field (run one)
        • M52 - JHK, dither 5' out of field - visible all night
        • M40 - Broad filters, few narrow filters - visible all night
    • G210 N1.8 LS Testing (Flexure and wavelength are completed)
      • Choose one
        • NGC 6052 - twilight-22hr
        • Arp 230 - 2nd half

PEPSI-PETS

Restart Continues

  • Resume above tasks depending on conditions and instrument availabiliy - DEPENDENT ON LUCI-1 MCE/ROE
  • If LUCI SW is fixed AND seeing < 1.2" AND no Field Aberration needed / AGW Transform data --> Proceed to AO (if Bin 3 and Bin 4 or ESM released)
    • Scripts: /home/lbto/Service_Observing/Restart_2022/LUCI/AO/AO_Testing_V94.xml
    • Full AO
      • Bin 3
      • Bin 4
    • ESM
      • Bin 1
      • Bin 2
      • Bin 4

LBC Chcekout

  • Skyflats if clear
  • Superfocus sequence
    • Requires 1.2" or better stable seeing
  • Photometric Standards
    • any seeing, must be clear
  • NS imaging?

Night Log

googly doc with info

-- BarryRothberg - 07 Sep 2022
Topic revision: r4 - 08 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback